OpenStudy (anonymous):

ITS A TRICKY PUZZLE Take 9 from 6, 10 from 9, 50 from 40 and leave 6. How Come ??

OpenStudy (anonymous):

SIX – 9 (IX)= S 9 (IX) – 10 (X) = I 40 (XL) – 50 (L)= X SIX

OpenStudy (anonymous):

using romans

OpenStudy (anonymous):

yeh, in romans only, i have gt dis

OpenStudy (anonymous):

ohh thats right dude @nitz

OpenStudy (anonymous):

thnx nitz

OpenStudy (anonymous):

another way: ((6-9) + (9- 10)) – (40-50)= -3 + -1 – (-10)= -4 + 10= 6

OpenStudy (anonymous):

any other puzzle?????

OpenStudy (anonymous):

yip that sounds abut right @nitz

OpenStudy (anonymous):

i will finding @nitz ..lolz

OpenStudy (anonymous):

i have got one puzzle

OpenStudy (anonymous):

which one?